Hi.I need your answer for this kuestion below.Br/H

A firm produces widgets according to the production function: F(K,L) = K 1/3L where K is capital and lLis labour. Does the firm’s production function exhibit constant- decreasing- or increasing returns to scale? Explain your answer.
Answer #1

Production function: Q = F(K, L) = K1/3L

When both inputs are doubled, new production function becomes

Q* = (2K)1/3(2L) = (2)1/3 x 2 x K1/3 x L = (2)4/3 x Q

Q*/Q = (2)4/3 > 2

Since doubling both inputs more than doubles output, there is increasing return to scale.